OpenStudy (anonymous):

how do you solve √8 + 3√2?

OpenStudy (lalaly):

√8=√4*√2 =2√2 so 2√2+3√2 = 5√2

OpenStudy (anonymous):

sqrt 8 = 2sqrt 2 so √8 + 3√2 = 2√2 + 3√2 = 5√2

OpenStudy (anonymous):

oh ok thankz